I'm confused and unsure of how to set up the following problem:
Ricky is 3 years older than his sister, Sara. 1 years ago, he was twice her age. How old is Ricky now?

Answer by stanbon(75887)

Ricky is 3 years older than his sister, Sara. 1 years ago, he was twice her age. How old is Ricky now?
-----------------
Equations:
r = s + 3
r-1 = 2(s-1)
----------
s = r-3
------------------
Substitute for s and solve for r::
r-1 = 2(r-3 -1)
r-1 = 2(r-4)
r-1 = 2r-8
r = 7 (Ricky's age now)
---------------
Cheers,
Stan H.
=============